# DexVLA: Vision-Language Model with Plug-In Diffusion Expert for General Robot Control

### v1

@@ IV Real World Experiments
IV-A Evaluating Model without Task-Specific Adaptation
This section evaluates the model’s performance before task-specific adaptation (Stage 3). The evaluated tasks, visualized in Figure 5, all use the model with one set of parameters.
@@ IV-A Evaluating Model without Task-Specific Adaptation
These tasks vary significantly in trajectory length and complexity, with some requiring high dexterity and intricate manipulation (e.g., shirt folding). We benchmark our approach against OpenVLA [23], a state-of-the-art 7B-parameter VLA model pre-trained on the Open X-Embodiment (OXE) [31] dataset, and Octo [30], a compact 93M parameter model that employs a diffusion-based policy for action generation. We used the open-sourced pre-trained weights for these two models. All baselines are fine-tuned on the same dataset and for the same number of epochs as our Stage 2 training, ensuring a fair comparison. We also compare to the Diffusion Policy [8], a strong baseline. Notably, neither Octo nor OpenVLA has previously demonstrated success on tasks of this complexity. While proprietary VLA models like $\pi_{0}$ and RT-2 [7] represent strong competitors, their closed-source implementations and reliance on private training data preclude direct empirical comparison.
Figure 8: Results on learning dexterous skills from new embodiment. We evaluated our model with four three baselines: Diffusion Policy, Octo, and OpenVLA. Diffusion Policy is directly trained on these novel tasks from scratch.
IV-B Learning Dexterous Skills on New Embodiment
This section evaluates our model’s ability to learn dexterous skills on the new embodiments as shown in Figure 7. Specifically, the new embodiments are not involved in either Stage 1 or Stage 2 training. We aim to demonstrate the effectiveness of our proposed framework in acquiring new skills quickly on any new embodiment without the necessity of pre-training.
@@ IV-B Learning Dexterous Skills on New Embodiment
These tasks involve two novel robotic systems absent from training data: 1) A Franka arm integrated with a dexterous hand have 12 DoF, serves as a more complex robotic system than a simple gripper, posing a greater challenge for the model to learn. 2) A bimanual UR5e system featuring humanoid-inspired kinematic design and its articulation fundamentally differs from conventional dual-arm platforms like the AgileX bimanual robot. We evaluate our approach against three baseline methods: Diffusion Policy, Octo, and OpenVLA, the same baseline as in the previous section. This section aims to validate the adaptability of our pre-trained model to new embodiments and tasks. To this end, we directly fine-tune our Stage 2 pre-trained model on the novel tasks. For OpenVLA and Octo, we employ their publicly available checkpoints pre-trained on the OXE dataset. The Diffusion Policy — a method specialized for learning dexterous tasks from limited data — is trained exclusively on the two novel tasks. All methods are trained on individual tasks, and to ensure fairness, each baseline undergoes the same number of training epochs.
IV-C Complex Long-Horizon Tasks with Direct Prompting
In this set of experiments, we tackle a range of challenging multi-stage tasks via a combination of task-specific post-training and self-generated reasoning capability. For some of these tasks, data is present in pre-training, but fine-tuning is required to attain mastery. For some, no data is present in pre-training. The tasks in this evaluation, shown in Figure 2, are:
@@ IV-C Complex Long-Horizon Tasks with Direct Prompting
These tasks involve extended-horizon challenges. For instance, laundry folding requires more than 2 minutes to collect a single episode, and the soft, deformable fabric of clothing generates numerous unseen shapes and states, posing significant challenges for recognition and task completion. In the sorting task, the model must pick up 5–8 randomly placed objects in a cluttered scene and relocate them to predefined target positions.
We evaluate all models using averaged scores over 10 trials (detailed scoring criteria are provided in the Appendix). Conduct comparisons on these tasks is challenging due to the scarcity of prior models capable of operating at this scale. While $\pi_{0}$ represents the state-of-the-art in laundry folding, as mentioned before, its closed-source implementation and reliance on proprietary training data prevent direct benchmarking. Consequently, we compare our method only against OpenVLA and Octo. As demonstrated in Figure 9, DexVLA consistently surpasses all baselines and achieves a 0.8 point in the dryer unloading task. Even facing the most complicated task, laundry folding, our method attains a 0.4 point which proves its potential in handling extremely complicated scenarios. Through comprehensive evaluations, DexVLA demonstrates the capability to perform versatile, complex, and dexterous tasks via the combination of our proposed embodied curriculum learning method, scalable diffusion expert, and our novel VLA framework. We argue that DexVLA provides a promising approach for constructing VLA systems that effectively handle heterogeneous robotics data and master intricate manipulation tasks.
V Conclusion
This work proposes DexVLA, a novel architecture that leverages vision-language models to learn semantic information and employs a billion-parameter diffusion expert to learn robust and generalizable visuomotor policies. We introduce an embodied curriculum learning strategy, enabling the network to progressively learn from embodiment-agnostic motor skills to complex, embodiment-specific dexterous skills through three training stages. Furthermore, we incorporate sub-step reasoning, allowing the model to perform very long-horizon tasks without relying on a high-level policy model. Our method is evaluated from multiple perspectives, including its ability to perform complex tasks without task-specific adaptation, fine-tune on new embodiments with limited data, and execute extremely complex, long-horizon tasks without the assistance of a high-level policy model. Across all tasks, our method outperforms state-of-the-art approaches such as OpenVLA, Octo, and Diffusion Policy. We also conduct ablation studies to verify the effectiveness and efficacy of our proposed method.
Limitations and future work. Our method has several limitations and areas for improvement. First, we did not observe a significant transfer of learning from data collected on other embodiments to a specific target embodiment. Second, we observed that complex tasks, such as laundry folding, strongly depend on accurate action recovery and the ability to recognize the correct object state within the image. Prior work, $\pi_{0}$, addresses this by using SayCan, where a high-level policy model frequently assesses the object state and provides updated language instructions to the low-level vision-language-action model. Our current method, relying solely on the vision-language component, struggles with this aspect. A potential solution is to design a mechanism for explicit state checking and integrate it into our end-to-end model, which we plan to investigate in future work. Similar to observations reported in $\pi_{0}$, we also found that models trained with stage 3 exhibit reduced cross-task generalization compared to models trained solely through stages 1 and 2. However, complex tasks like laundry folding necessitate stage 3 training to effectively learn the required actions and object states for successful completion. Addressing this trade-off between task-specific learning and generalization is crucial for enabling models to perform complex tasks effectively without relying on post-training.
VI Acknowledgement
We would like to thank Yanjie Ze for the valuable discussion. We also like to thank Xiaoyu Liu for contributing to data construction, and the efforts of Xiaoyu Liu, Zhongyi Zhou, Chengmeng Li, Yefei Chen, Yaxuan Li, and Peiting Zhao in data collection.

### v2

@@ 3.2 Embodied Curriculum Learning
Stage 3: Task-specific adaptation. This phase refines the model’s ability to execute downstream tasks skillfully and fluently, analogous to the post-training stage in large language models where the model is fine-tuned on domain-specific data. For simpler, less generalization-dependent tasks, such as shirt folding, table bussing, or bin picking with trained objects, task-specific training is unnecessary as the model already performs well. However, complex, dexterity-demanding tasks require the model to learn fine-grained, context-dependent actions. Therefore, effective post-training relies on a high-quality dataset of expert demonstrations exhibiting consistent and fluent task execution strategies focused on behaviors that promote successful task completion. It is worth noting that we utilize sub-step annotated language data in both Stage 2 and Stage 3. However, instead of directly using these sub-step reasoning as instructional input, we employ them as intermediate language output, compelling the model to learn and generate these sub-step language descriptions. This approach has proven highly effective, enabling our model to perform complex, long-horizon tasks such as laundry folding. While other VLAs, like $\pi_{0}$ [8], can also complete such tasks, they rely on high-level policy models like SayCan [35] to identify the task state and provide next-step instructions. In contrast, our framework leverages the VLM backbone as an implicit high-level policy model. This allows the model to internally interpret the task’s state and inject this understanding into the policy to guide action generation, eliminating the need for an external high-level policy module.
Substep reasoning. A key insight for training the DexVLA is the necessity of decomposing long-horizon tasks (e.g., bussing tables) into sub-tasks. These tasks, often spanning beyond 2 minutes, prove challenging for the diffusion expert to learn effectively from a single language instruction. Therefore, we annotate sub-step instructions within these long-horizon tasks to provide a more structured learning signal. Pre-training with sub-steps is crucial for strong performance. We demonstrate the importance of sub-step reasoning in the Appendix. Our empirical observations show that VLA trained without this pre-training frequently skips critical steps in very long tasks. Sub-step annotations are typically provided every five seconds of the demonstration. We show some examples for substep reasoning in Figure 2.
4 Real World Experiments
4.1 Evaluating Model without Task-Specific Adaptation
This section evaluates the model’s performance before task-specific adaptation (Stage 3). The evaluated tasks, visualized in Figure 6, all use the model with one set of parameters. Detailed task descrptions are listed in Appendix. These tasks vary significantly in trajectory length and complexity, with some requiring high dexterity and intricate manipulation (e.g., shirt folding). We benchmark our approach against OpenVLA [20], a 7B-parameter VLA model pre-trained on the Open X-Embodiment (OXE) [22] dataset, and Octo [21], a compact 93M parameter model that employs a diffusion-based policy for action generation. We used the open-sourced pre-trained weights for these two models. All baselines are fine-tuned on the same dataset and for the same number of epochs as our Stage 2 training, ensuring a fair comparison. We also compare to the Diffusion Policy [30], a strong baseline. Notably, neither Octo nor OpenVLA has previously demonstrated success on tasks of this complexity.
Figure 8: Results on learning dexterous skills from new embodiment. We evaluated our model with four baselines: Diffusion Policy, Octo, and OpenVLA. Diffusion Policy is directly trained on these novel tasks from scratch.
4.2 Learning Dexterous Skills on New Embodiment
This section evaluates our model’s ability to learn dexterous skills on the new embodiments as shown in Figure 8. Specifically, the new embodiments are not involved in either Stage 1 or Stage 2 training. We aim to demonstrate the effectiveness of our proposed framework in acquiring new skills quickly on any new embodiment without the necessity of pre-training. Detailed task descrptions are listed in Appendix.
These tasks involve two novel robotic systems absent from training data: 1) A Franka arm integrated with a dexterous hand has 12 DoF, serving as a more complex robotic system than a simple gripper, 2) A bimanual UR5e system featuring humanoid-inspired kinematic design and its articulation fundamentally differs from conventional dual-arm platforms like the AgileX bimanual robot. We evaluate our approach against the same baselines as in the previous section. This section aims to validate the adaptability of our pre-trained model to new embodiments and tasks. To this end, we directly fine-tune our Stage 2 pre-trained model on the novel tasks. For OpenVLA and Octo, we employ their publicly available checkpoints pre-trained on the OXE dataset. The Diffusion Policy — a method specialized for learning dexterous tasks from limited data — is trained exclusively on the two novel tasks. All methods are trained on individual tasks, and to ensure fairness, each baseline undergoes the same number of training epochs.
Figure 9: Average scores on tasks requiring stage 3 training. We compared our model against Octo, OpenVLA and $\pi_{0}$ on laundry folding and bussing table (hard).
Table 1: Ablation study of three stage training. To evaluate the impact of stage 1 training, we trained DexVLA from scratch without diffusion expert pre-training. Additionally, to assess the effectiveness of stage 3 training, we directly evaluated DexVLA on Laundry Folding without applying stage 3 training.
Stage 1 Stage 2 Stage 3 Shirt Folding Laundry Folding
- - 0.0 0.0
- - 0.0 0.0
- 0.92 0.0
0.92 0.4
Figure 10: Ablation results on size of Diffusion Expert. We reported the average score on the shirt folding task.
DexVLA UNet(93M) Diff. Expert(410M) Diff. Expert(1B)
Avg. Score 0.17 0.63 0.92
4.3 Complex Long-Horizon Tasks with Direct Prompting
In this set of experiments, we tackle a range of challenging multi-stage tasks via a combination of task-specific post-training and self-generated reasoning capability. For some of these tasks, data is present in pre-training, but fine-tuning is required to attain mastery. For some, no data is present in pre-training. The tasks in this evaluation, are shown in Figure 2, and detailed task descrptions are listed in Appendix. These tasks involve extended-horizon challenges. For instance, laundry folding requires more than 2 minutes to collect a single episode, and the soft, deformable fabric of clothing generates numerous unseen shapes and states, posing significant challenges for recognition and task completion. In the sorting task, the model must pick up 5–8 randomly placed objects in a cluttered scene and relocate them to predefined target positions. We report two tasks, laundry folding and bussing table (hard), in the main text, and discuss the rest of three tasks in the Appendix.
Our evaluation of all models is based on averaged scores over 10 trials, with detailed scoring criteria provided in the Appendix. Conducting comparisons on these tasks is challenging due to the limited availability of prior models capable of operating at this scale. Consequently, we compare our method specifically against OpenVLA, Octo, and $\pi_{0}$. Note that $\pi_{0}$ is pretrained on extensive data and shares the same embodiment as our tested robot. We use direct prompting on all models for fair comparison.
The results, illustrated in Figure 12, demonstrate that DexVLA consistently outperforms all baseline methods. For the most complex task—laundry folding—our method achieves a score of 0.4, showcasing its potential to handle highly complicated scenarios. In comparison, $\pi_{0}$ attains a score of 0.2 when directly prompted with task instructions. On the table bussing task, DexVLA surpasses $\pi_{0}$ by 0.08 points. These findings suggest that DexVLA presents a promising solution for executing complex, long-horizon tasks without relying on an external high-level planner.
4.4 Ablation Study
Ablation on three-stage training strategy. To validate the necessity of this multi-stage training process, we conduct an ablation study in this section. Specifically, we evaluate the DexVLA model under several training conditions: training only Stage 1, only Stage 2, both Stage 1 and Stage 2 combined, and all three stages. Stage 1 consists of pretraining using cross-embodiment data, Stage 2 involves fine-tuning with embodiment-specific data, and Stage 3 enables the model to master more complex tasks. The experimental results are summarized in Table 1. We observe that training solely on Stage 1 or Stage 2 leads to a 0% success rate in folding tasks. Notably, the absence of Stage 1 training results in the model completely failing to learn any meaningful actions. We hypothesize that the considerable number of parameters in the diffusion expert complicates the optimization process. Thus, Stage 1 serves not only to equip the diffusion expert with foundational action skills but also to ”warm up” its parameters, facilitating better comprehension of complex visual cues and language instructions. Additionally, to investigate the contribution of Stage 3 training to performance on more complex tasks, we directly evaluated DexVLA on the laundry-folding task without task-specific adaptation. We observed a significant performance drop from 0.4 to 0. These results highlight that Stage 3 training is essential for the model’s success in handling long-horizon and challenging tasks.
Ablation on size of diffusion expert. Our key contribution is a novel vision-language-action (VLA) model architecture incorporating a diffusion expert, a significantly larger action expert based on a diffusion transformer. However, does the larger diffusion transformer architecture of the diffusion expert (1B) offer advantages over a smaller one?
To address this question, we utilize a tiny UNet-based diffusion policy (93M) and a smaller diffusion expert (410M) as baselines. As shown in Table 10, the UNet-based action expert performs significantly worse than our method, barely completing the shirt folding task with an average score of 0.17. Empirically, we observed oscillation in the robot’s movements with the UNet model compared to our diffusion expert. We hypothesize that the UNet’s fewer parameters contribute to interference between different actions in the parameter space, hindering the model’s ability to learn the correct actions. The 410M diffusion expert achieved results lower than the 1B model (0.63 versus 0.92), indicating that learning a wide variety of tasks requires a greater number of model parameters.
5 Conclusion
This work proposes DexVLA, a novel architecture that leverages vision-language models to learn semantic information and employs a billion-parameter diffusion expert to learn robust and generalizable visuomotor policies. We introduce an embodied curriculum learning strategy, enabling the network to progressively learn from embodiment-agnostic motor skills to complex, embodiment-specific dexterous skills through three training stages. Furthermore, we incorporate sub-step reasoning, allowing the model to perform very long-horizon tasks without relying on a high-level policy model. Our method is evaluated from multiple perspectives, including its ability to perform complex tasks without task-specific adaptation, fine-tune on new embodiments with limited data, and execute extremely complex, long-horizon tasks without the assistance of a high-level policy model.
Appendix A More Experimental Results
A.1 Visual generalization.
Visual generalization is a critical aspect of robot learning. A well-trained model should not only perform well on in-domain tasks but also generalize to different objects within the same category and to novel scenes. This section presents our visual generalization tests. Specifically, we evaluate shirt folding on a bimanual AgileX and drink pouring using a Franka Emika robot with a dexterous hand. The former task is evaluated without task-specific adaptation, while the latter is trained with 100 demonstrations of the new embodiment. These tasks were also the focus of the experiments presented in Section 4.1 and Section 4.2, respectively. For both tasks, we assess visual generalization across two dimensions: novel objects and novel scenes.
Figure 11: Example of visual generalization. Here lists some visual generalization settings including unseen objects and unseen scenes.
Table 2: Visual Generalizaton for DexVLA. For each evaluation setting, we report the averaged scores across 3 trials.
Task / Generalization Embodiment Novel Object Novel Scene Novel Object & Scene
Shirt folding Bimanual AgileX 0.78 0.78 0.56
Drink pouring Dexterous hand 0.83 0.67 0.67
For shirt folding, we varied the shirt color (while maintaining size) and altered the background and scene. For drink pouring, we used unseen cups and bottles, also evaluating the task in different scenes and backgrounds. The results are presented in Table 2. Our experiments demonstrate that DexVLA effectively generalizes to novel visual environments. As shown in the supplementary video, the model successfully handles even challenging cases, such as folding white shirts on a white table. The examples are shown in Figure 11.
A.2 LIBERO simulation experimental results.
We compare our method on the Libero benchmark. DexVLA uses Stage 1 pre-trained weights. Specifically, the experimental results in Table 3 show that DexVLA outperforms all baselines, including $\pi_{0}$ and $\pi_{0}$-FAST, two state-of-the-art VLA methods, demonstrating strong performance on the benchmark.
Table 3: Evaluations results on LIBERO. We compare our DexVLA with DP and OpenVLA.
Method Spatial Object Goal Average
DP 78.3 92.5 68.3 79.7
OpenVLA 84.7 88.4 79.2 84.1
$\pi_{0}$-FAST 96.4 96.8 88.6 93.9
$\pi_{0}$ 96.8 98.8 95.8 97.1
DexVLA 97.2 99.1 95.6 97.3
A.3 Training cost of stage 1.
Table 4: Comparison of training cost for train only diffusion expert versus train entire VLA. Training cost is measured by the number of training epochs completed per hour.
Train Method Train only Diffusion Expert Train Entire VLA
Training Cost 0.89 epoch / hour 0.32 epoch / hour
As mentioned in Section 3.2, training the entire VLA model from scratch results in failure on nearly all tasks. Therefore, this section compares the training cost of our Stage 1 (training only the diffusion expert) with that of training the entire VLA model. The test reports the number of training epochs completed per hour. We deliberately keep the same batch size for fair comparison.
As shown in Table 4, training only the diffusion expert is 2.78 times faster than training the entire VLA model. This is expected, as the VLA model is three times larger than the diffusion expert alone. This highlights that our training strategy is not only effective but also cost-efficient.
A.4 More discussion on long-horizon tasks with direct prompting.
In 4.3, we previously compared DexVLA with $\pi_{0}$, Octo, and OpenVLA on laundry folding and bussing table hard. Here, we present additional results against OpenVLA and Octo on more complex tasks. As shown in Figure 12, DexVLA consistently outperforms all baselines, achieving 0.8 points in the dryer unloading task, while Octo and OpenVLA score 0 points. In tasks like sorting and bin picking hard, DexVLA achieves a score nearly 3 times higher than other baselines. Overall, DexVLA demonstrates its versatility in handling complex tasks through our embodied curriculum learning method, scalable diffusion expert, and novel VLA framework. We believe DexVLA presents a promising approach for building VLA systems capable of managing heterogeneous robotics data and mastering intricate manipulation tasks.
Figure 13: Robot setup on zero-shot cross-embodiment transfer from gripper to dexterous hand. We replace the original gripper with an Inspire dexterous hand and evaluate DexVLA in the same bin-picking environment using unseen objects.
A.5 Zero-Shot Cross-Embodiment Transfer
Finally, we pose an intriguing question: can DexVLA perform zero-shot cross-embodiment transfer? To explore this, we take a model trained on a simple two-finger gripper and deploy it, without any further training, on a dexterous five-finger hand. Specifically, we use the Stage 2 pre-trained DexVLA — which was trained on a Franka robot equipped with a Robotiq gripper — and swap in a dexterous hand at test time. Because the gripper has only one degree of freedom, we constrain the dexterous hand to a single degree of freedom as well.
We evaluate on a bin-picking task with 30 novel objects that were unseen during both Stage 1 and Stage 2 training. An illustrative example appears in Figure 13, and a full video demonstration is provided in the supplementary material. Across these 30 objects, we achieve an average success rate of 60%. While this is slightly below the 67% success rate attained with the original gripper, it nonetheless underscores the model’s robustness. The performance gap arises from three main challenges: (1) the dexterous hand’s appearance differs markedly from the gripper, forcing the model to generalize its visual feature representations; (2) the wrist camera’s mounting position shifts significantly, requiring adaptation to a new viewpoint; and (3) the difference in hand height changes the effective object grasping height. Although we do not yet control all degrees of freedom needed for fully dexterous manipulation, these results demonstrate that DexVLA’s visual and camera-view representations transfer effectively across embodiments.
Appendix B Ablation Study
B.1 Does training with substep reasoning help?
Table 5: Ablation study of substep reasoning. The ✓in each stage indicates the use of sub-step reasoning data during that stage. We report the average score on the shirt-folding task.
Stage 1 Stage 2 Averaged Score
0.07
0
0.92
A key strength of our method is its ability to handle extremely long and complex tasks, such as folding randomly crumpled shirts from a basket. It also enables the model to complete multi-stage tasks like shirt folding and bin picking without requiring post-training. Therefore, we now examine the importance of sub-step reasoning. We conducted an ablation study with two setups: 1) The diffusion expert is trained with direct prompting (each task has only one language instruction), while the VLA-diffusion expert is trained with sub-step reasoning. 2) Both stage 1 and stage 2 are trained with direct prompting data. The results are shown in Table 5. Training the diffusion expert with direct prompting, even for a relatively simple task like shirt folding, reduces the averaged score from 0.92 to 0.07. Furthermore, removing sub-step reasoning from both stages results in a complete failure (0 score). This is a significant observation. It suggests that learning long-horizon tasks within a shared parameter space can sometimes lead to conflicts. We hypothesize that sub-step reasoning allows the model to learn a more disentangled action space, similar to mapping a continuous action space to a discrete one [70]. This effectively segments the shared parameter space, allocating a smaller set of parameters to each substep [71, 72, 73]. This avoids parameter conflicts, leading to improved performance and generalization.
B.2 Explicit versus implicit sub-step reasoning.
Table 6: SayCan versus Substep Reasoning for DexVLA. We replace DexVLA’s substep reasoning with SayCan and evaluate how this change affects overall performance.
Tasks/Models SayCan Substep Reasoning
Bussing Table (Hard) 0.58 0.70
The $\pi_{0}$ utilizes the additional SayCan to explicitly plan substeps, while DexVLA generates substep reasoning implicitly. Here, we evaluate the impact of these two reasoning approaches on overall performance. Specifically, we replace DexVLA’s implicit substep reasoning with SayCan’s. As shown in Table 6, models using implicit substep reasoning significantly outperform those relying on SayCan. The key advantage of implicit substep reasoning is its ability to disentangle the action space across different features, allowing DexVLA to learn more effectively. Additionally, SayCan updates instructions at a fixed frequency (every two seconds), which can result in redundant or repeated states in certain scenarios. In contrast, DexVLA’s substep reasoning adaptively segments the state space over the course of long-horizon tasks, contributing to its superior performance.
Appendix C Task Suite and Evaluation Protocol
C.1 Task suite.
These tasks are evaluated in Section 4.1.
@@ Appendix C Task Suite and Evaluation Protocol
These tasks are evaluated in Section 4.2.
@@
These tasks are evaluated in Section 4.3.
@@
C.2 Evaluation protocol.
Each task is evaluated across 10 trials and reported averaged scores. For each task, we list the detailed scoring criterion as follows.
  • Lanudary folding (Bimanual AgileX): This task is scored out of 4 and we evaluate 5 shirts in total including 2 middle size and 3 small size. We perform two trials for each item, and the items left to be evaluated starting randomly crumpled in a laundry bin (while previously evaluated items start in a fold). One point is given for picking an item out of the bin and putting it on the table. Another point is given for flattening the shirt or shorts. A third point is granted for folding the shirt or shorts. A final point is given for either placing the item in the corner of the table (if it is the first item evaluated), or stacking it onto an existing stack of folded clothes. This evaluation metric is followed $\pi_{0}$.
  • Shirt folding (Bimanual AgileX): This task is scored out of 3. We perform two trials for each item, and the items are flattened on the table. One point is given for double vertically fold. Another point is granted for a double horizontal fold. A final point is given for pushing the folded shirt to the right blank area.
  • Bussing table (Bimanual AgileX): This task is scored out of 3-4 where there are 3-4 objects on the table in both easy and hard version. The main difference is the objects that appeared in the hard version are unseen. A point is given for each correctly sorted object.
  • Dryer unloading (Bimanual AgileX): This task is scored out of 2 where there are 2 crumpled shirts in the dryer. A point is given for pick up a shirt and place into the hamper.
  • Sorting (Franka with gripper): This task is scored out of 5-8 where there are 5-8 objects on the table. There are four kinds of objects in total, a point is given for each correctly sorted object.
  • Drink pouring (Franka with dexterous hand): This task is scored out of 2. A point is given for grab the bottle and pour to the cup. Another point is granted for place down the bottle.
  • Bining picking (Franka with gripper): This task is scored out of 4-5 where there are 4-5 objects on the table. The main difference is the objects that appeared in the hard version are unseen. A point is given for each correctly picked and placed object.
  • Packing (Bimanual UR5e): This task is scored out of 2 where there are 2 objects on the table. A point is given for each correctly picked and placed object.
Figure 14: Overview of our dataset for stage 1 training.
Figure 15: Average task length(seconds) of 91 tasks.
Appendix D More Implementation Details
D.1 Robot setup.
Our evaluation is conducted on four different robot configurations across 10 tasks. These setups are summarized in the following list and visualized in figure 4.
Our setup includes two Franka Emika robots: one equipped with a gripper and the other with a robot hand. Both Franka robots utilize the same camera configuration, consisting of a ZED 2 camera positioned on both the left and right sides, as well as a ZED Mini wrist camera mounted on the robot itself. Our bimanual UR5e robot uses a single top-mounted Intel RealSense L515 camera and two Intel RealSense 435i cameras attached to the wrists. Finally, our mobile AgileX platform has two Intel RealSense 435i wrist cameras and a top-mounted Intel RealSense 457 camera. Although the mobile AgileX image includes a front camera, it was not used during either training or inference.
D.2 Sub-step reasoning and data acquisition.
Training with sub-step reasoning is crucial for Dex-VLA to complete long-horizon tasks without a high-level policy model. We present an ablation study on the importance of sub-step reasoning in Section B.1. Acquiring this data presents two key challenges: obtaining language instructions and segmenting videos with corresponding annotations. We address these challenges with the following strategy.
For object-level tasks (e.g., bin picking, sorting, table bussing), object identification is key. We leverage Grounding-Dino and DINOv2 to annotate object bounding boxes and names, along with the gripper’s bounding box. We then calculate the intersection over union between the gripper and object bounding boxes to determine grasp success. For long-horizon single-object tasks (e.g., fold one shirt), the challenge lies in task segmentation. We created a comprehensive list of potential sub-step reasoning, focusing on major steps lasting at least five seconds each to avoid excessive sub-division. We then used Google Gemini 2.0, providing it with the sub-step list, to segment the videos and select the corresponding reasoning from the list. This proved effective and efficient for labeling. We only manually checked the Stage 3 training data, as this stage requires higher-quality annotations. This annotation strategy makes our approach feasible.
D.3 Architectural details.
In this section, we provide a full description of the model architecture. Dex-VLA can be split into two parts, VLM backbone originates from Qwen2-VL [27] and diffusion expert. We use Qwen2-VL 2B which is powerful and efficient. Regarding our Diffusion Expert, the total number of parameters for this model is 1 billion parameters. We use 32 layers, with the hidden stage of 1280, and a number of heads of 16. During Stage 1, we only pre-train the diffusion expert with random initialized ResNet-50 to process images and off-the-shelf Distilbert [31] to encode language instructions. Because the original diffusion policy model does not support cross-embodiment training, we adopted a multi-head structure similar to Octo [21]. Each embodiment is assigned a unique MLP head. The diffusion expert is trained using the similar settings of our Dex-VLA. In particular, we use the image resolution of 320 $\times$ 240, with three camera views. Each image is processed independently to a ResNet-50. We use the strategy as in RT-1 [33] to initialize the FiLM layers.
D.4 Training data details.
As shown in Figure 14. Our dataset comprises approximately 100 hours of collected data spanning 91 distinct tasks. The majority of this data was collected using two robot platforms: the Agilex (ARX arm) (42.7%) and the single Franka Emika robot (34.7%). The “ARX arm” and “PIPER arm” represent two distinct robotic arm configurations, both featuring six degrees of freedom (6-DoF) but differing in their kinematic structures and operational characteristics.
Figure 15 provides a summary of all 91 pretraining tasks across four embodiments. The Y-axis represents the average duration (in seconds) recorded for each task in the training data, while the X-axis lists all 91 tasks. It is clear that most tasks in the dataset are short-horizon, whereas the evaluated tasks are long-horizon, highlighting a notable distributional difference between the pretraining data and the evaluated tasks.